OpenStudy (anonymous):

4 - 2 × (-3 - 8) = ?

OpenStudy (rulnick):

26

Directrix (directrix):

26

OpenStudy (anonymous):

26

OpenStudy (rulnick):

Details: 4-2(-3-8) = 4-2(-11) = 4+22 = 26

OpenStudy (anonymous):

4 - 2 × (-3 - 8) = -11 * ( 2) = -22 4 -(- 22) = 26